Luebke Inc. has provided the following data for the month of November. The balance in the Finished Goods inventory account at the beginning of the month was $58,000 and at the end of the month was $30,600. The cost of goods manufactured for the month was $215,000. The actual manufacturing overhead cost incurred was $56,800 and the manufacturing overhead cost applied to Work in Process was $60,400. The company closes out any underapplied or overapplied manufacturing overhead to cost of goods sold. The adjusted cost of goods sold that would appear on the income statement for November is:

a) $238,800

b) 187,600

c) 242,400

d) 215,000
Answer #1

Answer a) $238800

Calculated as

Overhead over applied = $60400-56800 = $3600

So Cost of goods sold = $58000+215000-30600 = 242400

Adjusted cost of goods sold = $242400-3600 = $238800